2 ms·
Their answer to "Q: What's the probability that when we roll 2 d6s and add up the results, we get 7?" could be shorter. It is 1/6 because after the first die pr
by Bostonian 7y ago
Their answer to "Q: What's the probability that when we roll 2 d6s and add up the results, we get 7?" could be shorter. It is 1/6 because after the first die produces a number from 1 to 6, there is exactly a 1/6 chance that the second die will produce a number giving a sum of 7. The probability of getting a sum of 6 from two dice is 5/6*1/6 = 5/32 using similar logic.